the initial angular speed of the coil in the generator is 5 rad/s and the initial amplitude of the induced emf is 10 v.calculate the angular speed of the coil in the generator when the amplitude of the induced emf is 8v
1
Expert's answer
2019-09-30T10:16:32-0400

The emf induced in the coil of electrical generator is given by formula


"{\mathcal E}=NAB\omega\sin\omega t"

The amplitude of emf

"{\mathcal E_m}=NAB\omega"

Therefore


"\frac{\mathcal E_{m2}}{{\mathcal E_{m1}}}=\frac{\omega_2}{\omega_1}"

Finally

"{\omega_2}=\frac{\mathcal E_{m2}}{{\mathcal E_{m1}}}{\omega_1}"

"{\omega_2}=\frac{8\:\rm V}{10\:\rm V}\times{5\:\rm rad/s}=4\:\rm rad/s"